### v1

# Language steering in latent space to mitigate unintended code-switching

Abstract
Multilingual Large Language Models (LLMs) often exhibit unintended code-switching, reducing reliability in downstream tasks. We propose latent-space language steering, a lightweight inference-time method that identifies language directions via PCA on parallel translations and steers token embeddings along these axes to control language identity. Our approach mitigates code-switching while preserving semantics with negligible computational overhead and requires only minimal parallel data for calibration. Empirically, we achieve 95-99% language classification accuracy using a single principal component and reduce next-token distributional divergence by up to 42% across multiple language pairs on Qwen2.5 and Llama-3.2 models. We further analyze the layer-wise evolution of language representations, revealing that language identity concentrates in final layers with near-perfect linear separability. Code and data are released for reproducibility [^1] https://github.com/fxlrnrpt/language-steering-in-latent-space/tree/eacl2026.
Language steering in latent space to mitigate unintended code-switching
Andrey Goncharov1, Nikolai Kondusov, Alexey Zaytsev1, 1Applied AI Instiute Correspondence: andrey@faillearnrepeat.net
1 Introduction
Multilingual large language models (LLMs) frequently exhibit unintended code-switching - generating tokens in languages other than the target despite explicit monolingual instructions. This undermines reliability in production systems and complicates evaluation. Existing solutions require costly fine-tuning or introduce latency via decoding-time interventions like vocabulary masking or classifier guidance.
We propose latent space language steering, a cheap inference-time method that exploits low-dimensional language directions in hidden representations. By applying PCA to parallel translations, we identify linear subspaces separating languages while preserving semantics. Steering via simple projection (one dot product and vector subtraction per token) mitigates code-switching with negligible overhead.
Figure 1: Language steering in hidden space: 1. Collect parallel translations of the same concepts; 2. LLM forward pass to get embeddings by layer; 3. Perform PCA by layer on the merged embeddings.
Our analysis shows language identity concentrates in final layers, with PC1 capturing most of the variance and enabling near-perfect language classification. We validate across Qwen2.5 and Llama-3.2 models on multiple language pairs (English, Spanish, Russian, Chinese), demonstrating reduced distributional divergence while preserving fluency.
Contributions.
We characterize geometric structure of language identity in multilingual LLMs, introduce a cheap steering method with negligible overhead requiring only parallel translations for calibration, and empirically validate across models and language pairs, showing reduced code-switching while maintaining semantic fidelity.
2 Related works
Code-switching and multilingual control.
While code-switching has been studied for mixed-language inputs in speech and NER (Aguilar et al., 2020; Bali et al., 2014), recent work identifies unintended code-switching in multilingual LLM outputs (Ryan et al., 2024; Yoo et al., 2024), degrading user experience and evaluation. Standard mitigations require costly fine-tuning on monolingual data, reinforcement learning with language-specific rewards, or brittle prompt engineering.
Language structure in representations.
Work on mBERT and XLM-R showed language identity concentrates in few principal components, separable from semantics in parallel texts (Conneau et al., 2020; Chi et al., 2020). Yang et al. (Yang et al., 2021) removed these components for language-agnostic embeddings. Wendler et al. (Wendler et al., 2024) extended this to decoder-only LLMs, revealing latent language preferences. We exploit this structure for active generation control.
Inference-time control.
Plug-and-play methods steer models without fine-tuning via gradient-based attribute classifiers (Dathathri et al., 2020), future discriminators (Yang and Klein, 2021), or expert-guided decoding (Liu et al., 2021). These typically target sentiment or toxicity with auxiliary models or multiple forward passes. Our approach leverages language’s cleaner linear separability for simpler intervention: a single projection per token, no gradients, no auxiliary classifiers.
3 Methods
Our approach operates in the latent space of pre-trained multilingual LLMs without fine-tuning. We identify language-specific directions via PCA on parallel translations, then steer representations along these directions to control code-switching.
3.1 Language Direction Identification
Given a parallel corpus $\mathcal{D}=\{(s,\ell)\}_{i=1}^{N}$ where $s$ represents semantically equivalent content in language $\ell\in\mathcal{L}$, we extract hidden states $\mathbf{h}_{i}^{(\ell)}\in\mathbb{R}^{d}$ from each layer $\ell$ of the LLM. For each layer, we apply PCA to centered embeddings $\mathcal{H}^{(\ell)}=\{\mathbf{h}_{i}^{(\ell)}\}_{i=1}^{N}$. The first principal component $\mathbf{v}^{(\ell)}$ captures the language direction:
$\mathbf{v}^{(\ell)}=\arg\max_{\|\mathbf{v}\|=1}\sum_{i=1}^{N}\left(\mathbf{v}^{\top}(\mathbf{h}_{i}^{(\ell)}-\bar{\mathbf{h}}^{(\ell)})\right)^{2}$ (1)
When semantic content is constant, language identity becomes the dominant variance source, making $\mathbf{v}^{(\ell)}$ an interpretable language axis.
3.2 Latent-Space Language Steering
Empirical analysis (Section 4) shows language variance concentrates in final layers. We apply steering only to layers $\ell\geq\ell_{\text{crit}}$. For each token’s hidden state $\mathbf{h}_{t}^{(\ell)}$ during generation, we remove the language component:
$\tilde{\mathbf{h}}_{t}^{(\ell)}=\mathbf{h}_{t}^{(\ell)}-s\left(\mathbf{h}_{t}^{(\ell)}\cdot\mathbf{v}^{(\ell)}\right)\mathbf{v}^{(\ell)}$ (2)
where $s\in\mathbb{R}^{+}$ controls intervention strength. This removes language identity while preserving semantic content in the orthogonal subspace, adding only a dot product and vector subtraction per token.
3.3 Language Prediction
To validate our approach, we train a logistic regression classifier on final-layer projections $z_{i}=\mathbf{h}_{i}^{(L)}\cdot\mathbf{v}^{(L)}$:
$P(\ell\mid z)=\frac{\exp(w_{\ell}z+b_{\ell})}{\sum_{\ell^{\prime}\in\mathcal{L}}\exp(w_{\ell^{\prime}}z+b_{\ell^{\prime}})}$ (3)
This one-dimensional classifier predicts language identity with high accuracy, demonstrating the interpretability of discovered directions.
4 Results
(a) Final layer language map (Qwen)
(b) Final layer language map (Llama)
(a) First layer language map (Llama)
(b) Mid-layer language map (Llama)
(c) Explained variance (Llama)
(a) English - Chinese
(b) English - Spanish
(c) English - Russian
(d) English - Hindi
We validate our approach on Qwen2.5-1.5B  Qwen et al. (2025) and Llama-3.2-1B  Grattafiori et al. (2024) open models across multiple language pairs, demonstrating both the existence of interpretable language directions and their effectiveness for code-switching control.
4.1 Language direction identification
Experimental setup.
We use Flores Plus NLLB Team et al. (2024) dataset with 50 parallel samples (English, Spanish, Russian, Chinese, Hindi) for PCA fitting and 100 for validation.
Language maps reveal emergent clustering.
Figure 2(a) and 2(b) show final-layer embeddings projected onto the first two PCs. Despite no explicit language supervision, models spontaneously organize identical content into tight, language-specific clusters with geometrically precise boundaries. Similar language families (Spanish/Russian) cluster closer than typologically distant pairs (English/Chinese), suggesting PC1 captures coarse language distinction while PC2 encodes finer linguistic properties.
Figures  3(a)  3(b) show layer-wise evolution: clusters emerge loosely in early layers and sharpen dramatically in final layers. Figure 3(c) reveals late-stage specialization for language identity. Pairwise maps (Figures  4(a),  4(b),  4(c),  4(d)) confirm clean linear separability for all tested pairs.
4.2 Language prediction
Experimental setup.
We use Flores Plus NLLB Team et al. (2024) dataset with 50 parallel samples (English, Spanish, Russian, Chinese, Hindi) for PCA and logistic regression fitting and 100 for validation.
Classifier shows nearly perfect performance.
Table 1 demonstrates that a simple logistic regression classifier trained only on the first principal component achieves near-perfect accuracy (0.95-0.99) across all language pairs. This validates that language identity information is linearly accessible in the final layer representations, supporting our hypothesis that models maintain explicit language directions in their activation space. We can also clearly see the decreased performance for the English-Spanish pair. It can be attributed to the languages being much more closely related to each other.
Language pair Qwen
English - Chinese 0.98
English - Spanish 0.95
English - Russian 0.99
English - Hindi 0.98
Table 1: Last layer language classifier accuracy
4.3 Language steering
Experimental setup.
[…]
To find the best steering strength coefficient we conduct a grid search with the best values listed in Table 4 for each language pair.
Original Unsteered Steered
still muy very
going un surprised
not bastante pretty
so tan starting
a sor working
just más content
really a rather
in empez impressed
very de quite
sitting content
quite en great
pretty enc beginning
here aquí answering
thinking pens almost
looking realmente fully
kind dese $__$
getting completamente $_$
having feliz even
actually org heart
surprised seguro trying
Table 2: Top 20 token sample
Results.
Table 2 demonstrates that steering shifts predictions from Spanish tokens (muy, bastante, tan) to their English equivalents (very, quite, so) while preserving semantic coherence. It changing how the model speaks, not what it says.
Table 3 shows steering reduces divergence by 20% on average. It hints that while we can mitigate the code-switching, the complex structure of the manifold does not allow us to reconstruct the original concept with a simple linear transformation. At the same time, we see that the method shows superior performance for English-Chines language pair. We hypothesize that the root cause is the tokenization and training data imbalance. Most likely, Qwen model prcossed much more English and Chinese data compared to other language pairs. It leads to more granular tokenization and better internal generalization of the concepts.
Language pair Qwen
English - Chinese 8.94 $\rightarrow$ 5.19
English - Spanish 6.37 $\rightarrow$ 5.86
English - Russian 7.78 $\rightarrow$ 5.43
English - Hindi 8.73 $\rightarrow$ 8.73
Table 3: KL divergence (no steering $\rightarrow$ steering)
Language pair Qwen
English - Chinese -2.9
English - Spanish -1.4
English - Russian -2.2
English - Hindi 0
Table 4: Best steering strength coefficient
5 Conclusion
We introduced latent-space language steering, a lightweight inference-time method that mitigates unintended code-switching in multilingual LLMs without fine-tuning. By extracting language directions via PCA on parallel translations, our approach enables precise control over language identity with negligible overhead—one dot product and vector subtraction per token.
Our empirical analysis across Qwen2.5 and Llama-3.2 models reveals that language identity concentrates in final layers with near-perfect linear separability. This geometric structure enables both accurate language classification (95-99% accuracy) and effective steering that manages to restore the original language and reduces distributional divergence by up to 42% while maintaining semantic coherence. The layer-wise evolution shows language identity crystallizes in final layers while early layers focus on semantics, suggesting these occupy largely orthogonal subspaces.
While performance is strongest for typologically distant pairs (English-Chinese), the 20% average divergence reduction indicates that simple linear transformations cannot fully reconstruct monolingual distributions. Our approach demonstrates that interpretable control mechanisms can be extracted from learned representations, offering a practical alternative to expensive fine-tuning for production multilingual systems.
Limitations
  • Model scale and diversity. Experiments limited to small models (1-1.5B parameters) and two architectures (Qwen2.5, Llama-3.2). Larger models and diverse architectures remain unexplored.
  • Incomplete reconstruction. Average 20% KL divergence reduction indicates linear projections cannot fully reconstruct monolingual distributions, suggesting non-linear language-semantic interactions at the manifold level.
  • Dataset coverage. Evaluation relies on Flores Plus and TED Talks (formal text). Conversational data, technical domains, and naturally occurring code-switching patterns are not tested. Artificial code-switching setup may not reflect real-world scenarios.
  • Language typology. Limited to Indo-European and Chinese languages. Low-resource languages, right-to-left scripts, and typologically distant languages (agglutinative, tonal) unexplored. Hindi showed no improvement, suggesting tokenization or data balance issues.
  • Layer-wise steering. Steering applied only to final layers based on empirical observation. Systematic exploration of multi-layer or adaptive steering strategies not conducted.
  • Evaluation metrics. Primary reliance on KL divergence measures statistical similarity but not semantic preservation, fluency, or downstream task performance. Human evaluation and task-specific metrics needed.
  • Steering calibration. Optimal coefficients require manual grid search per language pair. Principled or adaptive selection methods not developed.
  • Efficiency benchmarking. No rigorous latency profiling across hardware configurations or batch sizes to validate negligible overhead claims for production deployment.